Consider an n-channel MOSFET (Von = 0.4 V and K = 3.0 mA/V2). Let VDD = 5.0 V, VSS = -5.0 V, R1 = 14.0 kohm, R2 = 6.0 kohm, RD = 1.2 kohm and RS = 0.5 kohm. Answer the following questions assuming the transistor is at its saturation mode.

a) Calculate VG versus ground (not VGS) (hint: voltage division by R1 and R2 between VDD and VSS).

b) Calculate VGS. (hint: IDS obtained by formula = IDS obtained by voltage drop between VG and VSS).

c) Calculate IDS.

d) Calculate VDS.

e) Verify the assumption of saturation mode using a calculation.
Answer #1

I SIDS Hizkni & luka PUTH 20.5k 20 VC 14... 84 6 Ta applying modal analysis at gate Von - VOD + VG =VSS=0 watt >> 2019 = vousJos = 3 (206 -0.550s) Iss = 3(6.776 7.0025 I35 -206 Ios) >> 6075175 – 8.8 Inst 20.2830 =) Tos= 3.13 mA, 8.6mA - Vs = 0.5(50s)